# Training-Free Activation Sparsity in Large Language Models

### v1

@@ 1 Introduction
This work describes TEAL (Training-Free Activation Sparsity in LLMs), a simple, training-free approach that applies activation sparsity based on magnitude pruning, where the threshold is layer-dependent. TEAL is based on the observation that distributional shapes in LLaMA-architecture LLMs are zero-mean unimodal. By pruning low-magnitude, non-salient activations, we achieve 40-50% model-wide (input-dependent) sparsity, in contrast to prior work which only achieves sparsity in portions of the model (Lee et al., 2024). We realize wall-clock speed-ups of up to 1.53×1.53$\times$1.53 × and 1.8×1.8$\times$1.8 × at 40% and 50% sparsity respectively through specialized kernels, and further demonstrate compatibility with weight quantization.
@@ 2 Related Work
Lee et al. (2024) propose CATS, and realize training-free activation sparsity on SwiGLU based LLMs by applying magnitude pruning on the output of $\textbf{W}_{\text{gate}}$, with the intuition that in the training-free setting, ReLU-based methods suboptimally zero out nontrivial negative values but keep positive values with lower magnitude intact. They achieve up to 50% sparsity in $\textbf{W}_{\text{up}}$ and $\textbf{W}_{\text{down}}$ for Mistral and Llama-2-7B. However, other matrices including $\textbf{W}_{\text{gate}}$ and $\textbf{W}_{\text{q,k,v,o}}$ are computed densely, resulting in lower model-wide sparsity (roughly 25%), whereas we target every matrix in the model.
@@ 4.1 Motivating Study: Distributional Properties of Activations in LLMs
We perform a preliminary study of the distributional properties of activations of LLMs. We collect activations of Llama-3-8B (Dubey et al., 2024) sampled from C4 (Dodge et al., 2021) at the four hidden states in a Transformer block, [^1] Throughout, we use “block” to refer to an entire Transformer layer consisting of the seven matrices and “layer” to refer to an individual layer within the Transformer block. and visualize them in Figure 2. As indicated by prior work, some of the activations are heavy-tailed and contain outliers (Dettmers et al., 2022; Xiao et al., 2022; Wei et al., 2022; Nrusimha et al., 2024). The hidden states are moreover zero-mean unimodal, and qualitatively fall into two distinctly shaped distributions. The hidden states before the Attention and the MLP layers tend to be Gaussian-like, while the hidden states in the intermediate of such layers tend to be Laplacian-like. The concentration of the activations around zero motivates our magnitude-based activation pruning approach.
@@ 4.3 Block-wise Greedy Optimization
For each Transformer block, we seek to minimize the block-wise $\ell_{2}$ activation error subject to a block-level sparsity constraint. Each Transformer block consists of seven matrices: $\mathbf{W}_{\text{q}},\mathbf{W}_{\text{k}},\mathbf{W}_{\text{v}},\mathbf{W}_{\text{o}},\mathbf{W}_{\text{gate}},\mathbf{W}_{\text{up}},\mathbf{W}_{\text{down}}$. Algorithm 1 initializes the sparsity levels of all layers to zero, and attempts to increment the sparsity level of each layer by an amount inversely proportional to its memory footprint. The layer with the lowest $\ell_{2}$ activation error is incremented, and the block-level sparsity plus associated layer-level sparsities are recorded. We share the block-level sparsity levels across all Transformer blocks; therefore, all blocks have the same target sparsity level, but the individual layer-level sparsities could be different across different blocks.
Cost. We describe the cost of our method. The time complexity is $\mathcal{O}(\frac{Mn^{2}}{\alpha})$ forward passes, where $M$ is the number of samples, $n$ is the number of matrices, and $\alpha$ is the average step size. In practice, we use length 2048 samples and $M,n,\alpha=10,7,0.05$. The resulting cost over all blocks is therefore $980016\cdot 7^{2}\cdot\frac{1}{0.05}=$ 9800 forward passes, which is less than one GPU-hour on an A100 for Llama-3-8B. It consumes minimal device memory due to its being block-wise and requiring no backpropagation.
@@ 4.4 Hardware Aware Acceleration
Figure 3 shows a small speed-up on A6000, and a larger speed-up on A100 over the DejaVu kernel. Note that torch.matmul is not the strongest baseline in small batch settings Hong et al. (2024), which is why our kernel is faster at 0% sparsity for A100. We use a stronger baseline for end-to-end evaluations (Section 5.2). The larger speed-up on A100 can be attributed to its higher memory bandwidth, which amplifies the impact of reducing other overhead factors. These overhead improvements become increasingly important as memory bandwidth across device tiers improves over time, particularly for quantized models and in latency-sensitive or resource-constrained applications.
@@ 5 Results
We use the block-wise greedily optimized sparsities from Section 4.3 for TEAL, and primarily compare to CATS (Lee et al., 2024) and ReLUfication (Mirzadeh et al., 2023) in their training-free configurations with no finetuning. We report model-level sparsities for all methods.
@@ 7 Conclusion
Acknowledgments
We thank Neil Movva, Jue Wang, and Yucheng Lu for helpful comments and discussion.

### v2

@@ 1 Introduction
This work describes TEAL (Training-Free Activation Sparsity in LLMs), a simple, training-free approach that applies activation sparsity based on magnitude pruning. TEAL is based on the observation that distributional shapes in LLaMA-architecture LLMs are zero-mean unimodal. By pruning low-magnitude, non-salient activations, we achieve 40-50% model-wide (input-dependent) sparsity, in contrast to prior work which only achieves sparsity in portions of the model (Lee et al., 2024b). We realize wall-clock speed-ups of up to 1.53×1.53$\times$1.53 × and 1.8×1.8$\times$1.8 × at 40% and 50% sparsity respectively through specialized kernels, and further demonstrate compatibility with weight quantization.
@@ 2 Related Work
Lee et al. (2024a) propose CATS, and realize training-free activation sparsity on SwiGLU based LLMs by applying magnitude pruning on the output of $\textbf{W}_{\text{gate}}$, with the intuition that in the training-free setting, ReLU-based methods suboptimally zero out nontrivial negative values but keep positive values with lower magnitude intact. They achieve up to 50% sparsity in $\textbf{W}_{\text{up}}$ and $\textbf{W}_{\text{down}}$ for Mistral and Llama-2-7B. However, other matrices including $\textbf{W}_{\text{gate}}$ and $\textbf{W}_{\text{q,k,v,o}}$ are computed densely, resulting in lower model-wide sparsity (roughly 25%), whereas we target every matrix in the model.
@@ 4.1 Motivating Study: Distributional Properties of Activations in LLMs
We perform a preliminary study of the distributional properties of activations of LLMs. We collect activations of Llama-3-8B (Dubey et al., 2024) sampled from C4 (Dodge et al., 2021) at the four hidden states in a Transformer block, [^1] Throughout, we use “block” to refer to an entire Transformer block consisting of the seven matrices and “layer” to refer to an individual layer (corresponding to a single matrix) within the Transformer block. and visualize them in Figure 2. As indicated by prior work, some of the activations are heavy-tailed and contain outliers (Dettmers et al., 2022; Xiao et al., 2022; Wei et al., 2022; Nrusimha et al., 2024). The hidden states are moreover zero-mean unimodal, and qualitatively fall into two distinctly shaped distributions. The hidden states before the Attention and the MLP layers tend to be Gaussian-like, while the hidden states in the intermediate of such layers tend to be Laplacian-like. The concentration of the activations around zero motivates our magnitude-based activation pruning approach.
@@ 4.3 Block-wise Greedy Optimization
For each Transformer block, we seek to minimize the block-wise $\ell_{2}$ activation error subject to a block-level sparsity constraint. Each Transformer block consists of seven matrices: $\mathbf{W}_{\text{q}},\mathbf{W}_{\text{k}},\mathbf{W}_{\text{v}},\mathbf{W}_{\text{o}},\mathbf{W}_{\text{gate}},\mathbf{W}_{\text{up}},\mathbf{W}_{\text{down}}$. Algorithm 1 initializes the sparsity levels of all layers to zero, and attempts to increment the sparsity level of each layer by an amount inversely proportional to its memory footprint. The layer with the lowest $\ell_{2}$ activation error is incremented, and the block-level sparsity plus associated layer-level sparsities are recorded. We assign the same block-level sparsity level across all Transformer blocks; therefore, all blocks have the same target sparsity level, but the individual layer-level sparsities could be different across different blocks.
Cost. We describe the cost of our method. The time complexity is $\mathcal{O}(\frac{Mn^{2}}{\alpha})$ forward passes, where $M$ is the number of samples, $n$ is the number of matrices, and $\alpha$ is the average step size. In practice, we use length 2048 samples and $M,n,\alpha=10,7,0.05$. The resulting cost over all blocks is therefore $980010\cdot 7^{2}\cdot\frac{1}{0.05}=$ 9800 forward passes, which is less than one GPU-hour on an A100 for Llama-3-8B. It consumes minimal device memory due to its being block-wise and requiring no backpropagation.
@@ 4.4 Hardware Aware Acceleration
Figure 3 shows a small speed-up on A6000, and a larger speed-up on A100 over the DejaVu kernel. Note that torch.matmul is not the strongest baseline in small batch settings (Hong et al., 2024), which is why our kernel is faster at 0% sparsity for A100. We use a stronger baseline for end-to-end evaluations (Section 5.2). The larger speed-up on A100 can be attributed to its higher memory bandwidth, which amplifies the impact of reducing other overhead factors. These overhead improvements become increasingly important as memory bandwidth across device tiers improves over time, particularly for quantized models and in latency-sensitive or resource-constrained applications.
@@ 5 Results
We use the block-wise greedily optimized sparsities from Section 4.3 for TEAL, and primarily compare to CATS (Lee et al., 2024a) and ReLUfication (Mirzadeh et al., 2023) in their training-free configurations with no finetuning. We report model-level sparsities for all methods.